# Risk premium and rough volatility

### v1

@@ 3.1. Estimating the risk premium in rough Bergomi
$\left\{\begin{array}[]{rl}\displaystyle\frac{\mathrm{d}S_{t}}{S_{t}}&=\displaystyle\mu_{t}\mathrm{d}t+\sqrt{v_{t}}\,\mathrm{d}W^{\mathbb{P}}_{t},\\ v_{t}&=\displaystyle\exp\left\{\nu Z^{H}_{t}\right\},\\ \end{array}\right.\quad\text{vs}\quad\left\{\begin{array}[]{rl}\displaystyle\frac{\mathrm{d}S_{t}}{S_{t}}&=\displaystyle r_{t}\mathrm{d}t+\sqrt{v_{t}}\,\mathrm{d}W^{\mathbb{Q}}_{t},\\ v_{t}&=\displaystyle\xi_{0}(t)\exp\left\{\nu\left(\int_{0}^{t}\mathtt{k}_{H_{-}}(t-s)\mathrm{d}Z^{\mathbb{Q}}_{s}+\int_{0}^{t}\mathtt{k}_{H_{-}}(t-s)\lambda_{s}\mathrm{d}s\right)\right\},\end{array}\right.$
@@
The deterministic process $\lambda$ admits the following piecewise constant representation on the time partition $\{0=T_{0}<T_{1},...,<T_{n}=T\}$, namely for $n\in\mathbb{N}$:
(3.10) $\lambda(t):=\sum_{i=1}^{n}\lambda_{i}1{1}_{\{t\in[T_{i-1},T_{i})\}},\quad\lambda_{i}\in\mathbb{R}\text{ for }i=1,\ldots,n.$
Similarly the forward variance $\xi_{0}$ admits the following piecewise constant representation $\mathbb{E}_{0}^{\mathbb{Q}}[v_{t}]=\xi_{0}(t):=\sum_{i=1}^{n}\xi_{i}1{1}_{\{t\in[T_{i-1},T_{i})\}}$ with $\xi_{i}\in\mathbb{R}$ for $i=1,\ldots,n$, where $\xi_{i}:=\frac{\mathfrak{V}_{T_{i}}T_{i}-\mathfrak{V}_{T_{i-1}}T_{i-1}}{T_{i}-T_{i-1}}$ and $\mathfrak{V}_{T}:=\mathbb{E}^{\mathbb{Q}}\left[\frac{1}{T}\int_{0}^{T}v_{s}\mathrm{d}s\right]$ is a market variance swap quote.
[…]
@@
(3.11) Parameters $\text{Parameters }H,\nu,\rho,\qquad\mathbb{E}^{\mathbb{P}}[v_{t}|\mathcal{F}_{0}],\qquad\mathbb{E}^{\mathbb{Q}}[v_{t}|\mathcal{F}_{0}].$
So far we have obtained $\mathbb{E}^{\mathbb{Q}}[v_{t}|\mathcal{F}_{0}]$ from Variance Swap market quotes. The next step is to estimate $(H,\nu,\rho)$ using historical time-series. Gatheral, Jaisson and Rosenbaum [10], explain how to estimate $\widehat{H}$ and $\widehat{\nu}$ from daily volatility data (Figure 3), and we follow their approach using a 100-day rolling window (Figure 4) and refer the reader to the original paper for details.
@@
In order to estimate the correlation parameter we use the formula
(3.12) $\mathrm{Corr}\left(Z^{H}_{t}-Z^{H}_{s},\int_{s}^{t}\mathrm{d}W_{s}\right)=\frac{\rho\sqrt{2H}}{H_{+}},$
which allows us to estimate the correlation with the proxy
@@
To forecast volatility and obtain $\mathbb{E}^{\mathbb{P}}[v_{t}|\mathcal{F}_{0}]$, we proceed as in [10] and use the forecasting formula for the fractional Brownian motion due to Nuzman and Poor [22]:

### v2

@@ 3.1. Estimating the risk premium in rough Bergomi
[…]
@@
Given a time partition $\{0=T_{0}<T_{1},...,<T_{n}=T\}$, the deterministic process $\lambda$ admits the piecewise constant representation
(3.8) $\lambda(t):=\sum_{i=1}^{n}\lambda_{i}1\hskip-5.97508pt{1}_{\{t\in[T_{i-1},T_{i})\}},\quad\lambda_{i}\in\mathbb{R}\text{ for }i=1,\ldots,n.$
Similarly the forward variance admits the piecewise constant representation $\mathbb{E}_{0}^{\mathbb{Q}}[v_{t}]=\xi_{0}(t):=\sum_{i=1}^{n}\xi_{i}1\hskip-5.97508pt{1}_{\{t\in[T_{i-1},T_{i})\}}$ with $\xi_{i}\in\mathbb{R}$ for $i=1,\ldots,n$, where $\xi_{i}:=\frac{\mathfrak{V}_{T_{i}}T_{i}-\mathfrak{V}_{T_{i-1}}T_{i-1}}{T_{i}-T_{i-1}}$ and $\mathfrak{V}_{T}:=\mathbb{E}^{\mathbb{Q}}\left[\frac{1}{T}\int_{0}^{T}v_{s}\mathrm{d}s\right]$ is a market variance swap quote.
We now estimate $\{\lambda_{1},\cdots,\lambda_{n}\}$. The dataset consists of daily Eurostoxx variance swap quotes for maturities $\{\text{1M, 3M, 6M, 1Y, 2Y}\}$ (Figures 1 and 2), while Figure 3 shows the daily realised volatility obtained from Oxford-Man institute data.
@@
(3.9) $\text{Parameters }H,\nu,\rho,\qquad\mathbb{E}^{\mathbb{P}}[v_{t}|\mathcal{F}_{0}],\qquad\mathbb{E}^{\mathbb{Q}}[v_{t}|\mathcal{F}_{0}].$
So far we have obtained $\mathbb{E}^{\mathbb{Q}}[v_{t}|\mathcal{F}_{0}]$ from Variance Swap market quotes. The next step is to estimate $(H,\nu,\rho)$ using historical time-series. Gatheral, Jaisson and Rosenbaum [gatheral2022volatility], explain how to estimate $\widehat{H}$ and $\widehat{\nu}$ from daily volatility data (Figure 3), and we follow their approach using a 100-day rolling window (Figure 4) and refer the reader to the original paper for details.
@@
To estimate the correlation parameter we use $\mathrm{Corr}\left(Z^{H}_{t}-Z^{H}_{s},\int_{s}^{t}\mathrm{d}W_{s}\right)=\frac{\rho\sqrt{2H}}{H_{+}}$, which allows us to estimate the correlation with the proxy
@@
To forecast volatility and obtain $\mathbb{E}^{\mathbb{P}}[v_{t}|\mathcal{F}_{0}]$, we proceed as in [gatheral2022volatility] and use the forecasting formula for the fractional Brownian motion due to Nuzman and Poor [nuzman2000linear]:
@@
Appendix A Proof of Remark 1.5
Following the ideas in [gassiat2019martingale, Proof of Theorem 1.1], we show that Assumption 1.3(iv) is guaranteed provided that Assumption 1.1 and 1.3(i)-(ii) hold and the processes $\mu,\gamma,r$ belong to $\mathbb{F}_{b}$. From the definition of the change of measure and the stopping time in (1.6), then $\mathcal{D}^{\gamma}_{t\wedge\tau_{n}}=\frac{\mathrm{d}\widehat{\mathbb{P}}_{n}}{\mathrm{d}\mathbb{P}}\bigg|_{\mathcal{F}_{t}}$. For any $s\in\mathbb{T}$, the random function $f(x):=\frac{r_{s}-\mu_{s}}{\sqrt{\psi(s,x)}}$ is $\mathbb{P}$-bounded on $(-\infty,a]$ for any $a>0$ since $r$ and $\mu$ are $\mathbb{P}$-bounded, and $\psi(s,\cdot)$ bounded away from zero on intervals of the form $(-\infty,a]$, by Assumption 1.3(i) together with the additional assumptions in Remark 1.5. Then, again by Proposition 1.2,
(A.1) $1=\mathbb{E}\left[\mathcal{D}^{\gamma}_{T\wedge\tau_{n}}\right]=\mathbb{E}\left[\mathcal{D}^{\gamma}_{T}1\hskip-5.97508pt{1}_{\{T<\tau_{n}\}}\right]+\mathbb{E}\left[\mathcal{D}^{\gamma}_{\tau_{n}}1\hskip-5.97508pt{1}_{\{\tau_{n}\leq T\}}\right].$
The first term in (A.1) converges to $\mathbb{E}\left[\mathcal{D}^{\gamma}_{T}\right]$ as $n$ tends to infinity, yielding
$1-\mathbb{E}\left[\mathcal{D}^{\gamma}_{T}\right]=\lim_{n\uparrow\infty}\mathbb{E}\left[\mathcal{D}^{\gamma}_{\tau_{n}}1\hskip-5.97508pt{1}_{\{\tau_{n}\leq T\}}\right].$
Girsanov’s theorem implies $\mathbb{E}\left[\mathcal{D}^{\gamma}_{\tau_{n}}1\hskip-5.97508pt{1}_{\{\tau_{n}\leq T\}}\right]=\widehat{\mathbb{P}}_{n}(\tau_{n}\leq T)$, where $\widehat{\mathbb{P}}_{n}$ is defined such that $\widehat{W}^{n}_{t}=W^{\mathbb{P}}_{t}-\int_{0}^{t\wedge\tau_{n}}\chi_{u}\mathrm{d}u$ is a $\widehat{\mathbb{P}}_{n}$-Brownian motion. Then, under $\widehat{\mathbb{P}}_{n}$, the process $Y$ becomes
$Y_{t}=\widehat{Y}_{t}^{n}+\int_{0}^{t\wedge\tau_{n}}k(t,s)\Big(\rho\chi_{u}+\overline{\rho}\;\gamma_{u}\Big)\mathrm{d}u=\widehat{Y}_{t}^{n}+\int_{0}^{t\wedge\tau_{n}}k(t,s)\lambda_{u}\mathrm{d}u,$
where $\widehat{Y}^{n}_{t}:=\int_{0}^{t}k(t,s)\mathrm{d}\widehat{Z}^{n}_{s}$ and
$\widehat{Z}^{n}_{t}=Z^{\mathbb{P}}_{t}-\int_{0}^{t\wedge\tau_{n}}\Big(\rho\chi_{u}+\overline{\rho}\;\gamma_{u}\Big)\mathrm{d}u=Z^{\mathbb{P}}_{t}-\int_{0}^{t\wedge\tau_{n}}\lambda_{u}\mathrm{d}u,$
for $t\geq 0$, where $\widehat{Z}^{n}$ is a $\widehat{\mathbb{P}}_{n}$-Brownian motion. Furthermore, by Assumption 1.3(ii), we have
(A.2) $\widehat{\mathbb{P}}_{n}\left(\sup_{t\in\mathbb{T}}Y_{t}\geq n\right)$ $=\widehat{\mathbb{P}}_{n}\left(\sup_{t\in\mathbb{T}}\left\{\widehat{Y}_{t}^{n}+\int_{0}^{t\wedge\tau_{n}}k(t,u)\lambda_{u}\mathrm{d}u\right\}\geq n\right)$
$\leq\widehat{\mathbb{P}}_{n}\left(\sup_{t\in\mathbb{T}}\widehat{Y}_{t}^{n}+\sup_{t\in\mathbb{T}}\left\{\int_{0}^{t\wedge\tau_{n}}k(t,u)\lambda_{u}\mathrm{d}u\right\}\geq n\right)$
$\leq\widehat{\mathbb{P}}_{n}\left(\sup_{t\in\mathbb{T}}\widehat{Y}_{t}^{n}\geq n-K_{\mathbb{T}}\right),$
Inequality (A.2), in turn, implies $\widehat{\mathbb{P}}_{n}(\tau_{n}\leq T)\leq\widehat{\mathbb{P}}_{n}(\widehat{\tau}_{n}\leq T)$, for $\widehat{\tau}_{n}:=\inf\{t\geq 0,\;\widehat{Y}_{t}^{n}=n-K_{\mathbb{T}}\}$. Finally, since $\widehat{Z}^{n}$ is a $\widehat{\mathbb{P}}_{n}$-Brownian motion, we obtain
$\lim_{n\uparrow\infty}\widehat{\mathbb{P}}_{n}(\tau_{n}\leq T)\leq\lim_{n\uparrow\infty}\widehat{\mathbb{P}}_{n}(\widehat{\tau}_{n}\leq T)=\lim_{n\uparrow\infty}\mathbb{P}\left(\sup_{t\in\mathbb{T}}Y_{t}\geq n-K_{\mathbb{T}}\right)=0,$
and it follows that $\frac{\mathrm{d}\mathbb{Q}}{\mathrm{d}\mathbb{P}}$ is indeed a true martingale and note that $\lim_{n\uparrow\infty}\widehat{\mathbb{P}}_{n}=\mathbb{Q}$. In the sense that the relation (1.8) holds between the $\mathbb{P}$ and $\mathbb{Q}$ Brownian motions.
Licence and Data Access statement
For the purpose of open access, the author(s) has applied a Creative Commons Attribution (CC BY) licence (where permitted by UKRI, ‘Open Government Licence’ or ‘Creative Commons Attribution No-derivatives (CC BY-ND) licence’ may be stated instead) to any Author Accepted Manuscript version arising’
The data underpinning this study were obtained from a combination of publicly and commercially available sources:
  • Oxford-Man Institute of Quantitative Finance Realized Library: Available for academic research use at https://realized.oxford-man.ox.ac.uk/ upon registration and agreement to the terms of use.
  • Yahoo Finance: Market data were retrieved from https://finance.yahoo.com, a publicly accessible platform, under their data usage terms.
  • Bloomberg Terminal: Data sourced from Bloomberg are subject to licensing restrictions and cannot be shared publicly. Access to Bloomberg data requires an institutional subscription.
Some components of the dataset are restricted due to commercial licensing agreements, and therefore cannot be made openly available. The decision to restrict access is to comply with the contractual obligations agreed with commercial data providers. The present authors may be contacted for further information about the datasets used here.